OpenStudy (anonymous):

if (x) = 2x^3+Ax^2+4x-5 f(2)=5 what is the value of A?

OpenStudy (experimentx):

f(x) = 2x^3+Ax^2+4x-5 then f(2) = ?

OpenStudy (anonymous):

5

OpenStudy (experimentx):

if it is not given then? what would you do?

OpenStudy (anonymous):

I dont know but it is given here . very new to concepts of function. plz help me solve this equation

OpenStudy (experimentx):

f(x) = 2x^3+Ax^2+4x-5 f(a) = 2a^3+Aa^2+4a-5 f(b) = 2b^3+Ab^2+4b-5 f(2) = ?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

f(2)=5

OpenStudy (experimentx):

f(2) = 2(2)^3+A(2)^2+4(2)-5

OpenStudy (anonymous):

so u plug in the values of x=2

OpenStudy (experimentx):

yes .. and f(2) = 5 2(2)^3+A(2)^2+4(2)-5 = 5

OpenStudy (anonymous):

ok so 16+A4-8-5?

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

A4-3

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=Solve [2%282%29^3%2BA%282%29^2%2B4%282%29-5+%3D+5%2CA]

OpenStudy (experimentx):

4A - 3 = 5 A = (5+3)/4 = 2